### v1

# Ten Years of ZMap

Abstract
Since ZMap’s debut in 2013, networking and security researchers have used the open-source scanner to write hundreds of research papers that study Internet behavior. In addition, ZMap powers much of the attack-surface management and security ratings industries, and more than a dozen security companies have built products on top of ZMap. Behind the scenes, much of ZMap’s behavior—ranging from its pseudorandom IP generation to its packet construction—has quietly evolved as we have learned more about how to scan the Internet. In this work, we quantify ZMap’s adoption over the ten years since its release, describe its modern behavior (and the measurements that motivated those changes), and offer lessons from releasing and maintaining ZMap.
1. Introduction
In 2013, Durumeric et al. released ZMap (durumeric2013zmap, ), an open-source network scanner that made it dramatically easier to scan the entire IPv4 address space. Since then, more than 300 research papers have used ZMap to uncover protocol flaws (adrian2015imperfect, ; aviram2016drown, ), shed light on the WebPKI, reverse engineer mercenary spyware (marczak2014governments, ), understand headline events like Heartbleed (durumeric2014matter, ) and Mirai (antonakakis2017understanding, ), and more. Beyond research, more than a dozen security companies have developed products on top of ZMap to continuously monitor companies’ attack surfaces and their third-party dependencies. At the same time, like many security tools, ZMap has also been adopted by attackers to identify vulnerable systems. In aggregate, ZMap accounts for over one-third of all Internet scan traffic today.
Since its initial release, ZMap has greatly evolved as we have learned more about how to scan the Internet. Yet, with the exception of Adrian et al.’s write up of ZMap’s 10GbE re-architecture (adrian2014zippier, ), the project team has not documented these changes in the research literature. Indeed, of the improvements Adrian et al. introduced, both the lock-free randomization algorithm and fast packet transmission approach have changed. We have also learned a great deal about how to (and not to) build measurement tools.
Motivated by requests from the community to document what we learned (claffy2019workshop, ), in this work, we present a retrospective analysis of ZMap. We cover how ZMap has been adopted (Section 2), advances in Internet scanning that challenged the tool’s initial design assumptions (Section 3), significant changes to ZMap’s behavior (Section 4), and what we learned maintaining and improving ZMap (Section 5). We hope that by analyzing this endeavor, we can help the researchers who create the next generation of Internet measurement tools.
2. Usage and Internet Landscape
Since ZMap’s release in 2013, its adoption has steadily increased (durumeric2014internet, ; anand2023aggressive, ). Today, over 33% of all IPv4 scan traffic can be fingerprinted as coming from ZMap. In this section, we describe ZMap’s adoption by academic researchers, security companies, and malicious actors.
Figure 1. ZMap-Attributed Scan Traffic—ZMap growth has accelerated significantly since 2020. In Q1 2024, 35% of global IPv4 TCP scan traffic (by packet) came from ZMap.
2.1. Empirical Analysis
One year after ZMap’s release, in 2014, Durumeric et al. found that there had been relatively little adoption and the tool was primarily being used to study academically interesting protocols like HTTP(S) (durumeric2014internet, ). More recently, in 2021, Anand et al. noted that many “aggressive” scans were from ZMap or Masscan (anand2023aggressive, ). Using the same methodology as these two studies (durumeric2014internet, ; anand2023aggressive, ), we measure ZMap adoption by analyzing scans that target at least 10 IPs in the ORION Network Telescope (orion, ). Our analysis is limited to TCP scans, since ORION does not tag scanning tools for ICMP and UDP flows.
ZMap usage has increased dramatically over the past four years (Figure 1). From January 1 to March 31, 2024, 35.4% of all IPv4 TCP scan packets originated from ZMap. While it is possible that ICMP and UDP scans use other tools, the vast majority of all scan traffic is TCP-based and ZMap still accounts for 33% of all scan packets with only TCP traffic attributed. ZMap scans follow a different distribution of ports than global scans (Figures 2 and 3). For example, ZMap accounts for only 12% of TCP/23 (Telnet) but 69% of TCP/80 and 73% of TCP/8080. In the most extreme case, 99.5% of traffic to TCP/8728 (MikroTik router API) is from ZMap, driving it to the sixth most scanned port globally. There are also dramatic regional differences in ZMap adoption (Figure 4). For example, while ZMap accounts for more than 66% of scan packets from U.S. hosts, less than 0.5% of Russian scan packets are from ZMap. The outsized use in the U.S. is largely due to adoption by American security companies (Section 2.3).
Figure 2. All TCP Scans (Top 30 Ports by Packet)
Figure 3. Top Ports Scanned by ZMap (by Packet)
2.2. Academic Research
ZMap has been used for a vast range of research purposes, from compromising RSA keys through transient faults  (sullivan2022open, ) to measuring NAT64 deployment (hsu2024first, ). To understand the studies that ZMap has enabled, we manually analyzed 1,034 papers that cite or reference ZMap through April 2024 via Google Scholar. We exclude dissertations from our analysis (since these are often comprised of published papers) as well as studies that use Censys (durumeric2015search, ). In total, we identified 307 research papers directly based on ZMap data.
While ZMap is a general measurement tool, it has most prominently been used by the security community. Notably, ZMap has been used in 38 papers to uncover protocol weaknesses in TLS and underlying cryptographic primitives (beurdouche2015messy, ; adrian2015imperfect, ; aviram2016drown, ) as well as to uncover deployment challenges and measure adoption (nemec2017measuring, ; kranch2015upgrading, ; holz2015tls, ). Collecting X.509 certificates through scans, 28 papers have shed light on the WebPKI prior to the widespread adoption of Certificate Transparency (vandersloot2016towards, ). There is also a large contingent of papers that have measured the Internet exposure of IoT devices (25 papers), ICS (14 papers), and security-relevant services (12 papers). Beyond understanding deployment patterns, a number of papers have been able to identify real-world attacks (durumeric2015neither, ), to reverse engineering attacker infrastructure (antonakakis2017understanding, ; marczak2015pay, ; marczak2014governments, ), and to conduct large-scale security notifications using ZMap data (li2016exploring, ; durumeric2014matter, ).
More networking-focused papers cover topics like DNS (24 papers), BGP/RPKI (12 papers), censorship (14 papers), and IP usage and NAT (10 papers). In addition to these studies, 53 other papers reference following our recommended practices when conducting other measurements. We provide a comprehensive breakdown of topics in Appendix B.
US NL RU DE GB BG CN IN ZA HK
66% 33% 0.48% 18% 69% 9% 2% 12% 0.1% 2%
Figure 4. ZMap by Country—The ten countries that emanate the most scan traffic by packet have varied ZMap usage.
2.3. Industry Adoption
Despite the plethora of publications, academic institutions are not responsible for most ZMap traffic. None of the top 100 ASes that emit the most ZMap traffic belong to universities; rather, most traffic originates from security companies and cloud providers. The network responsible for by far the most traffic is Google Cloud, which we find is predominantly used by Palo Alto Networks to power their Xpanse product.
To understand industry adoption, we categorized the organizations identified by Greynoise as using ZMap. We identify several categories of security products that are built on ZMap and fast Internet-wide scanning:
Attack Surface Management. With the shift to cloud-based infrastructure and a rise of ransomware attacks against enterprise services (e.g., MoveIt and VMWare ESXi), there has been an increased demand for companies to understand their Internet-facing infrastructure. Palo Alto Xpanse, Microsoft RiskIQ, and Rapid7 insightVM, along with more than a dozen other smaller companies, use ZMap as the basis for providing “attack surface management” products that give enterprises up-to-date data about their Internet-exposed risks and potentially unknown assets (“Shadow IT”).
Third-Party Risk Management. Building on the observation that externally visible security configuration and patching patterns correlate with data breaches and compromise (liu2015cloudy, ; zhang2014mismanagement, ; liu2015predicting, ), companies such as BitSight and FICO use ZMap to build security ratings that enable companies to understand the security of their supply-chain dependencies. SecurityScorecard, another major vendor, uses Masscan (ssc-masscan, ).
Internet Intelligence. More generically, several non-profit and commercial organizations use ZMap to collect data about IP addresses, threat actors, and Internet services, including BinaryEdge, Censys, IPInfo, and ShadowServer. Through these data sources, multiple countries now proactively monitor and notify organizations about risks (e.g., U.K. (ncsc-censys, )).
2.4. Malicious Use
Most security tools have the potential for both helping defenders and being misused by attackers. ZMap is no exception and there is evidence that ZMap has been used maliciously. While a significant proportion of scan traffic comes from networks with strong reputations (e.g., Google Cloud), attackers have also used “bulletproof” hosting providers to carry out scans for vulnerable services, including MSSQL, RDP, and Mikrotik’s router API (durumeric2014internet, ; anand2023aggressive, ). Attackers also have begun to repurpose ZMap to carry out DOS attacks (russia-dos, ), and, recently, two IoT botnets incorporated ZMap into their malware: in 2021, Palo Alto Networks found a Mirai variant adopted ZMap to search for SSH services (mirai-zmap, ), and, in 2023, a Medusa variant adopted ZMap to find Telnet services (medusa-zmap, ).
3. Lessons in Internet Scanning
Subsequent discoveries about Internet scanning have challenged some of ZMap’s original design assumptions:
Port Diffusion. Despite IANA assignment of ports to L7 protocols, Bano et al. (bano2018scanning, ) first noted and Izhikevich et al. (izhikevich2021lzr, ) more formally showed that protocols run across a long tail of ports: only 3% of HTTP and 6% of TLS services run on ports 80 and 443, respectively (izhikevich2021lzr, ). Scanning only assigned ports works well for understanding common user-facing protocols such as HTTP(S) but underestimates the impact of some security phenomenon, such as malware and industrial control system exposure. This has spurred new research into more intelligent Internet scanning approaches (sarabi2021smart, ; izhikevich2022predicting, ; song2023doors, ; luo2024ipreds, ) and led us to shift ZMap’s address generation from being purely “horizontal” to support multiple ports.
L4 vs. L7 Discrepancies. Several studies have noted significant discrepancies between L4 and L7 responsiveness (mirian2016internet, ; durumeric2013analysis, ; holz2015tls, ; springall2016ftp, ; heninger2012mining, ; perino2018proxytorrent, ). Izhikevich et al. showed that TCP liveness does not reliably indicate the presence of an application-layer service because of pervasive middlebox deployment (izhikevich2021lzr, ). ZMap’s design also encouraged what Hiesgen et al. term “two-phase scanning” (hiesgen2022spoki, ), in which L4 service discovery and L7 service interrogation are performed separately. In response, some hosts “shun” two-phase scanners, exacerbating the perceived differences between L4 and L7 results (izhikevich2021lzr, ). Sattler et al. later devised a method for more accurately identifying highly L4-but-not-L7 responsive prefixes (sattler2023packed, ). These differences fundamentally limit ZMap’s utility (as a standalone L4 tool) to discovering potential services, requiring most work to be completed in follow-up L7 scans and shifting our focus to downstream tools like LZR (izhikevich2021lzr, ) and ZGrab (zgrab2, ).
Visibility and Consistency. One challenge of running Internet scans is lack of ground truth for validation. Wan et al. showed that the ZMap paper slightly overestimated the visibility achieved by the tool, and that a single-probe scan actually misses about 2.7% of HTTP(S) hosts (wan2020origin, ). As can be seen in Figures 1 of Hastings et al. (hastings2016weak, ) and Chung et al. (chung2016measuring, ), different organizations using ZMap sometimes see different results. However, in even the most egregious cases (e.g., Censys (durumeric2015search, )), vantage points only miss under 5% of services due to blocking (in addition to transient drop); the bulk of loss is typically driven by a handful of small service and cloud providers (wan2020origin, ).
We now recommend that the best way to mitigate transient drop is to scan from 2–3 geographically and topologically diverse vantages, rather than to send multiple probes from a single scanner, since both probes are oftentimes lost. Results can also differ across scanning tools: Adrian et al. showed that, despite following a similar high-level approach, Masscan (graham2014masscan, ) finds notably fewer hosts than ZMap, likely due to biases in Masscan’s randomization algorithm (adrian2014zippier, ).
4. ZMap Codebase
When we released ZMap, we had little idea what community, if any, would emerge. Over the past ten years, more than 80 individuals have committed code to ZMap—we are deeply thankful to those contributors. Despite the high number of committers, 90% of ZMap code has been written by five individuals. Most contributions to the code have been made by industry researchers and software developers. Academic involvement has been limited: of the 11 external contributors who changed more than 100 LoC, only one is an academic researcher. Academic groups most frequently contribute probe modules or bugfixes; when academic groups made improvements to core functionality, improvements tend to be forked and renamed (e.g., XMap (xmap, ) and ZMapv6 (gasser2016scanning, ), which implemented the same IPv6 functionality) rather than upstreamed.
Beyond usability improvements and bug fixes, there have been several significant changes in how ZMap operates that make the original paper’s description of the tool out of date. We describe the most significant of these changes below.
4.1. Address and Port Generation
[…]
Hosts vs. Targets. While ZMap originally tracked statistics by IP address, our randomization approach selects from a pool of (IP, port) “targets”, rather than considering IPs and ports independently. As a result, options like “max hosts” cannot be supported. All results, statistics, and configuration in ZMap are now based on the notion of an IP/port target.
Identifying Generators. To create a new permutation of the address space for every scan, ZMap originally identified a random generator (i.e., primitive root) of the appropriate multiplicative group by identifying a generator of ($\mathbb{Z}_{p-1}$, +) and then mapping it into ($\mathbb{Z}^{*}_{p}$, $\times$)  (durumeric2013zmap, ). This was practical because a generator of the additive group is any integer coprime with $p-1$. This can be efficiently tested against randomly drawn integers, and there are $\phi(2^{32}+14)\approx 10^{9}$ generators of the additive group, resulting in an average of four attempts to identify a generator. Nearly all additive generators could be mapped into usable generators in the multiplicative group, since the only constraint on the multiplicative generators was that they were less than $2^{32}$ (to ensure safe multiplication using 64-bit arithmetic).
For multiport scans, to support iterating over $2^{48}$ elements, we need to efficiently find multiplicative generators smaller than $2^{16}$ in a $2^{48}$ search space. Unfortunately, because an additive generator can map to a multiplicative generator anywhere in the group, only $1/2^{32}$ candidate additive generators are usable. To address this, we flipped our approach. For each group defined by prime $p$, we precalculate and store the prime factorization of $p-1=k_{1}^{a_{1}}k_{2}^{a_{2}}\dots k_{n}^{a_{n}}$ for distinct primes $k_{i}\,(1\leq i\leq n)$. At runtime, we generate a random candidate generator $g\in[2\mathrel{{.}\,{.}}\nobreak 2^{16}-1]$ that is guaranteed to keep future arithmetic within the 64-bit address space. Then we ensure that $g$ is a generator of $\mathbb{Z}^{*}_{p}$ by checking that: $g^{(p-1)/k_{i}}\mod p\neq 1\,\forall\,i\in[1\mathrel{{.}\,{.}}\nobreak n]$. Similarly to the first approach, this requires an average of four attempts.
Figure 5. Sliding Window Duplicate Rate—We moved to a sliding window-based approach for deduplicating responses as part of supporting multiple ports. A window size of at least $10^{7}$ eliminates nearly all duplicates.
Response Deduplication. Hosts frequently send back repeated responses, in some cases indefinitely. While we initially thought these were due to broken TCP implementations, Goldblatt et al. showed that some hosts will aggressively send tens of thousands of response packets in response to a SYN packet, which they term “blowback” (goldblatt2023blowback, ). We originally filtered out duplicate responses using a paged $2^{32}$-bit bitmap, which used 512 MB of memory. While this approach guarantees no duplicates, extending it to the 48-bit space of IPs and ports would require 35 TB. Instead, we switch to maintaining a sliding window of the last $n$ IP/Port responses, using a Judy array (baskins2000judy, ). As can be seen in Figure 5, maintaining a window of $10^{7}$ entries is effective to filter nearly all duplicate responses, and lower scan rates can make do with smaller window sizes. We found zero duplicates for the three trials of 1 Gbps scans targeting TCP/80 in April 2024.
4.2. Sharding
In 2014, Adrian et al. (adrian2014zippier, ) introduced a mutex-free sharding mechanism that enabled spreading address generation across multiple senders without synchronization. This enabled a single Internet-wide scan to be split across multiple machines and improved performance by allowing multiple send threads on one machine to operate independently. As Mazel et al. noted when they showed that ZMap can be fingerprinted through its IP generation method (mazel2019identifying, ), we further shifted the sharding approach in 2017, as described below.
(a) Interleaved (Old)
(b) Pizza (New)
Figure 6. Sharding Approaches—In 2017, ZMap changed sharding approaches. Replicated with permission from (mazel2019identifying, ).
Interleaved Sharding. Sharding was originally implemented with each shard iterating by $N$ steps at a time with offset starting elements. For example, for $N$ shards, each shard $n\in[1\mathrel{{.}\,{.}}N]$ iterates by multiplying the current element by $g^{N}$, but begins iteration at $g^{n}$ (Figure 6(a)). With multiple threads in place, each shard $n$ is then further split into $T$ subshards, with each subshard iterating by $g^{NT}$ offset by $g^{n+tN}$, where $n$ is the shard index and $t$ is the thread index.
While conceptually simple, the approach requires calculating the end point of each shard to know when to stop iterating. $NT$ is not guaranteed to cleanly divide $p-1$, and so a shard might not repeat its initial value until it finishes iterating the full group (rather than just the shard). Unfortunately, the last index of a shard does not have a closed form expression, and we found that calculating it is prone to off-by-one errors. After repeated such correctness issues, we switched to a simpler sharding mechanism.
Pizza Sharding. Rather than interleaving shards, we divide the multiplicative group into $N$ ranges of values of increasing order, e.g., $[g^{0},g^{(p-1)/N})$, $[g^{(p-1)/N}$, $g^{2(p-1)/N})$, $[g^{2(p-1)/N}$, $g^{3(p-1)/N})$. For subshards, we similarly slice a single shard into $T$ ranges of values of increasing order. Visually, this is similar to slicing a pizza into $N$ slices, and then subdividing each slice into $T$ subslices (Figure 6(b)). Because elements are iterated over pseudorandomly in the group, the same randomness guarantees are provided by the second approach while being easier to reason about and implement without off-by-one errors or infinite loops.
4.3. Packet Construction
Striving for the highest send rate, ZMap originally used the smallest possible probes, with no included TCP or IP options. While protocol compliant, we later observed that ZMap would consistently miss some hosts accessible to OS network stacks. By varying TCP options, we found that including any of the Selective ACK (SA), Timestamp (TS), Window Select (WS), or Maximum Segment Size (MSS) TCP options yields a 1.5–2.0% increase in hitrate relative to no options in a scan of TCP/80 (Figure 7). The order of options also affects results: the optimal byte-layout order, minding the TCP 4-byte word boundary, finds only slightly fewer hosts (0.0023%, $\approx$1.5K hosts in an IPv4 scan of TCP/80) than when sending options using the exact ordering of Linux, BSD, or Windows.
TCP options affect packet size and therefore scan rate. However, including the MSS option alone finds the vast majority of services (over 99.99% of services on TCP/80) and remains under the minimum size of an Ethernet frame, continuing to support the maximum 1.488 Mpps line rate of a 1 GbE link. Using the Windows or Linux packet layouts finds very slightly more services but reduces send rates to 1.389 and 1.276 Mpps, respectively. While filtering packets based on TCP options could be due to defensive mechanisms attempting to block scanning, removing the easily identified static IP ID value of ZMap probes appears to have no impact on scan hit rates. We performed three scans of 10% of IPv4 on TCP/80 in April 2024 with a static IP ID and with a random per-packet IP ID and find that the difference in hit-rate between the random and static IP IDs is not statistically significant. In early 2024, ZMap changed to using random per-probe IP IDs.
Figure 7. Hitrates for Varying TCP Options—TCP SYN probes without any TCP options, as originally sent by ZMap, find 1.5–2.0% fewer services on TCP/80 than including any options. Mimicking common OSes maximizes coverage.
5. Lessons and Recommendations
Based on our experiences maintaining ZMap and feedback we received from other users, we offer the following lessons and recommendations for future measurement tools.
Tools Not Frameworks. First documented in 1978 by McIlroy et al. (mcilroy1978unix, ) and more concisely captured by Salus in 1994 (salus1994quarter, ), the Unix philosophy is to write programs that:
  1. (1)
    do one thing and do it well;
  2. (2)
    work together; and
  3. (3)
    handle text streams (a universal interface).
Nearly 50 years later, we cannot agree more with this guidance. It is impossible to fully predict how researchers will use measurement tools or the environments where they will operate (e.g., running ZMap on 16-bit embedded systems). ZMap was originally envisioned as a framework where researchers would build customized Scan Modules and Output Modules for service follow up. In practice, the vast majority of researchers use ZMap for service discovery and pipe results to secondary tools for investigation or storage. The output modules ZMap included for connecting to specific databases (e.g., Redis) became liabilities, requiring upkeep and complicating testing and packaging. In time, we removed them, opting to support only Text, CSV, and JSON Lines output.
Recommendation: Build small, simple, easy-to-understand, easy-to-use, and easy-to-test measurement tools that can be creatively assembled, rather than complex applications or software frameworks. Build applications that do one thing well. Continuously output results on a per-record/per-line basis when possible. Avoid proprietary formats and standardize output on well-worn interfaces like JSON Lines (jsonlines, ) or CSV, even if those formats are less computationally efficient; interoperability with other tools is more important.
Usability. ZMap was not the first fast, asynchronous network scanner: Unicornscan (unicornscan, ), Scanrand (scanrand, ), and IRLscanner (leonard2010demystifying, ) were released years prior (though Unicornscan and Scanrand were both unknown to the originating ZMap team at the time). ZMap was more successful than these tools due to its greater ease of use: it enabled researchers to scan the Internet from a single machine by running a single command.
Recommendation: Obsess over ease of installation, usage, and troubleshooting. It is better to have a tool that is easy to use and less full featured than vice versa.
Library and Command Line Wrapper. It is natural to build an application where the command-line interface, application configuration, and operation are intermingled, since most tools are first used via the command line and this is the path of least resistance. However, this design will limit a tool’s potential to be integrated into larger systems. Automated or continous measurements, such as the scans that power services like Censys (durumeric2015search, ), are cumbersome to control from the command line and more suited to a library interface.
Recommendation: Structure tools with two major components: a backend library and a simple command line interface that wraps the library. This investment is relatively small and will enable the tool to be used in larger systems.
Data, Metadata, and Logs. Unix has two standard output streams: Standard Output (stdout) and Standard Error (stderr). We argue that, in contrast, measurement tools should produce three types of output: (1) Data [Standard Output], (2) Logs [Standard Error], and (3) Metadata. Given the amount of data produced, it is difficult to tell whether an experiment is operating as expected without analyzing metadata in real time. In addition, tracking as much information about the execution (e.g., time, version of software, configuration parameters, and environment) will help to later interpret, troubleshoot, or reproduce results. Logs are helpful for human debugging, but they are not inherently designed to be machine-parsable, which is needed for monitoring long-running experiments. Ultimately, we extended ZMap to produce four output streams: (1) data, (2) logs, (3) real-time updates (e.g., packets sent, received, dropped per second), and (4) machine-readable metadata at completion.
Recommendation: Design measurement tools to produce separate streams of data, metadata, and debug logs. Do not mix these streams, since this complicates downstream processing. Be liberal in what environment and execution information is included in scan metadata, as it is difficult to know a priori what will be useful. Adopt a logging library that supports multiple log levels, and use debug-level logging liberally to enable future troubleshooting.
Static Types and Output Schema. JSON and CSV provide considerable flexibility when encoding data. For example, in JSON, objects can have dynamic keys, and values are not typed. However, downstream applications and databases usually do not support this flexibility, and it is easy to create valid JSON records that are painful to process.
Recommendation: Define a schema for the data you output. Ensure that each field uses a single, well defined type and that the type of one field does not depend on the value of another field. Avoid maps with dynamic keys, and instead use lists of a static document type. Consider using a tool like JSON Schema (json-schema, ) or ZSchema (zschema, ) to document the structure of output data.
Versioning and Releases. We released ZMap far too infrequently, repeatedly falling victim to wanting to include one more feature or bugfix in each release. In particular, ZMap 3.0 was released nearly eight years after the previous release of ZMap 2.1.1. Unfortunately, this meant that most users were either using long out-of-date releases or running HEAD. This made debugging difficult and prevented users from describing which ZMap version they used.
Recommendation: Follow the Semantic Versioning Specification (semver, ) religiously. Focus on making regular versioned releases rather than trying to finish a preset amount of work.
Language Choice. When we wrote ZMap, C and C++ were the only practical high-performance systems programming languages available. It is deceivingly easy to implement software in C and to convince oneself that it is possible to safely write C code. Empirical evidence overwhelmingly says the opposite (gaynor2020science, ), and even the White House now recommends moving towards memory safe languages (whitehouse2024memory, ). Network parsers are particularly hard to implement safely and must be resistant to attacker-controlled and malformed data (chromeruleoftwo, ). Over the years, ZMap has had multiple regressions that caused incorrect or incomplete measurements, as well as multiple memory safety bugs, all of which could have been avoided with a different choice of languages. Our team has also found that memory safety concerns make it harder to review external contributions for safety and correctness, which has reduced the rate at which we can merge improvements.
Recommendation: Responsibly develop measurement tools in modern, memory-safe languages like Go and Rust. While Rust has a relatively steep learning curve, its safety and runtime performance make it ideal for performance-critical applications. Go’s performance, while less than Rust’s, is sufficient for many measurement applications (e.g., ZGrab (zgrab2, ) and ZDNS (zdns, )), and the language’s simple syntax and parallelism-oriented architecture make it particularly suited for quickly developing network measurement tools.
6. Conclusion
The most exciting aspect of building ZMap has been watching how other researchers have used it in a myriad of unpredicted but valuable ways. We are sincerely thankful to everyone who has contributed, pushing the tool close to being feature complete. While the development of new ZMap features has slowed, we are excited to continue to expand the ecosystem of tools that work with ZMap (e.g., ZDNS (zdns, ) and ZGrab (zgrab2, )) and thereby to enable an even broader range of measurement uses. Many of the lessons we learned from maintaining ZMap may seem obvious today, but we did not have the foresight originally. We hope that our retrospective analysis will help the community build an even richer and more reliable ecosystem of Internet measurement tools moving forward.
Appendix A Ethics
Our work is primarily a metareview of prior studies and a presentation of lessons learned from ZMap. When conducting our own experiments to validate changes made to ZMap, we followed the original guidelines set forth by Durumeric et al. in 2013 (durumeric2013zmap, ). However, revisiting the release of ZMap, this work also provides us with an opportunity to revisit the original recommendations for how best to conduct active Internet-wide scans (durumeric2013zmap, ):
  1. (1)
    Coordinate closely with local network admins to reduce risks and handle inquiries.
  2. (2)
    Verify that scans will not overwhelm the local network or upstream provider.
  3. (3)
    Signal the benign nature of the scans in web pages and DNS entries of the source addresses.
  4. (4)
    Clearly explain the purpose and scope of the scans in all communications.
  5. (5)
    Provide a simple means of opting out, and honor requests promptly.
  6. (6)
    Conduct scans no larger or more frequent than is necessary for research objectives.
  7. (7)
    Spread scan traffic over time or source addresses when feasible.
Overall, we believe that these recommendations remain a sound set of considerations when conducting active Internet measurement. However, we also encourage the Internet measurement community to treat these recommendations as a set of good practices for most research, not as a set of requirements nor as the basis for having conducted research ethically. For example, there may be some situations where scans may be better performed unattributed, or where opting out certain networks could invalidate an operation (e.g., when tracking a specific threat actor). Institutions have adopted different practices for validating opt-out requests . Given that IP addresses ownership can change, it may make sense to expire opt-out requests.
Beyond these original recommendations, we additionally recommend that researchers consider the following:
  1. (1)
    Investigate whether existing datasets (e.g., Censys Research Access and Rapid7 Open Data) can suffice. In many cases, existing datasets can provide higher coverage of services across a large number of ports than researcher-operated scans while simultaneously reducing the aggregate bandwidth used by scans.
  2. (2)
    Publish newly collected datasets for other researchers.
  3. (3)
    Deploy WHOIS entries for research subnets that help operators to directly contact researchers.
  4. (4)
    Validate how handshakes will appear in the logs of scanned systems. For example, some SSH handshakes that do not attempt to authenticate still show up as failed authentication attempts causing undue worry for operators monitoring their logs.
We offer the following, updated set of best practices:
  1. (1)
    Minimize Internet Impact. While Internet scanning is a powerful methodology, it can also affect systems and create work for operators. Consider whether existing open source datasets provide the data you need. If you do perform scans, conduct scans no larger or more frequent than is necessary for research objectives. Publish data for future researchers.
  2. (2)
    Signal Intent. When possible, publish reverse DNS entries, IP WHOIS records, and an informational website that describes the benign nature of scans and how to contact the responsible researchers.
  3. (3)
    Provide An Opt-Out Mechanism. Provide a simple mechanism for operators to request to be opted out of future scans. Indicate the IP ranges you use for scanning so that operators can drop research traffic.
  4. (4)
    Proactively Investigate Effects. Run any newly developed scanning code against your own systems to ensure that you understand how scans might affect devices and appear in system logs. Start with small experiments before completing full Internet scans in case your scanner causes unexpected problems.
  5. (5)
    Coordinate Locally. Coordinate with your local IT and security teams to reduce the risk of overwhelming local networks, as well as to ensure that they know how to handle any inbound inquiries.
  6. (6)
    Disclose Results. When conducting security studies, consider how you can improve the security of the systems you have scanned. Responsibly disclose security problems you find and consider notifying individual system owners who are vulnerable.
Appendix B Academic Usage of ZMap
Topic Papers Examples
Censorship and Anonymity 14 (ramesh2020decentralized, ; pearce2017augur, ; pearce2017global, ; frolov2020detecting, )
Cryptography and Key Generation 17 (bos2014elliptic, ; janovsky2020biased, ; vsvenda2016million, ; hastings2016weak, )
Denial of Service (DoS) 15 (giotsas2017inferring, ; bushart2018dns, ; bock2021weaponizing, )
DNS and Naming 24 (liu2016all, ; lu2019end, ; zhu2015measuring, )
Email and Spam 8 (szurdi2017email, ; maroofi2021adoption, ; durumeric2015neither, ; holz2015tls, )
Exposure, Hygiene, and Patching 12 (bonkoski2013illuminating, ; fiebig2016one, ; zhang2014mismanagement, ; durumeric2014matter, )
Honeypots, Telescopes, and Attacks 9 (franzen2022looking, ; alt2014uncovering, ; morishita2019detect, )
IP Usage, DHCP Churn, Nand AT 10 (moura2015dynamic, ; lee2016identifying, ; hsu2024first, )
Industrial Control Systems (ICS) 14 (mirian2016internet, ; dahlmanns2020easing, ; dahlmanns2022missed, )
Internet of Things (IoT) 25 (mangino2020internet, ; antonakakis2017understanding, ; costin2014large, ; srinivasa2021open, )
Systems and Network Security 19 (xu2015measurement, ; vissers2015maneuvering, ; guo2018abusing, ; malhotra2015attacking, )
PKI, Certificates, Revocation 28 (liu2015end, ; brubaker2014using, ; durumeric2013analysis, ; perl2014you, )
Power Outages and Grid Monitoring 4 (bayat2021down, ; andersonpowerping, )
Privacy 5 (perta2014exploiting, ; van2022saving, ; hilts2015half, )
QUIC 7 (ruth2018first, ; zirngibl2021s, ; kosek2022dns, )
Routing, BGP, and RPKI 12 (giotsas2015mapping, ; alaraj2023global, ; rodday2021deployment, ; hlavacek2020disco, )
Scanning and Device Identification 25 (sattler2023packed, ; yang2019towards, ; albakour2021third, ; albakour2023pushing, )
TLS, HTTPS, and SSH 38 (adrian2015imperfect, ; brinkmann2021alpaca, ; holz2015tls, ; springall2016measuring, )
Understanding Threat Actors 4 (marczak2014governments, ; marczak2015pay, )
Other Internet Measurement Topics 26 (mehani2015early, ; rye2019sundials, ; guo2018detecting, ; lichtblau2017detection, )
Ethics Guidance Only (No ZMap Use) 53 (nguyen2022freely, ; kim2018measuring, ; izhikevich2024democratizing, ; moon2021accurately, )
Figure 8. Academic Papers Built on ZMap Data—We manually investigated papers that cited ZMap or ZMap-derived datasets to understand what types of research studies have used ZMap data.